It has charged a accumulator with an unknown electrical capacity with a electric potential difference of 99.73 V.

And when the accumulator was connected in parallel to a 10.0 μF electrical capacity or to another uncharged accumulator,

the electric potential difference in potential for each accumulator was 32.0 V.

What is the electrical capacity of the first accumulator?

p.s I cant understand too. but it is only.
Answer #1

Q = CV V = 99-73N V = 32 V al = 32C cl2l0uF (@-@) sen at al = 0 a (Ac) a ( + ) - a ct ce a 32 c² + c. 32c a ac 32c +32c = 2